# Nearly-Linear Time Seeded Extractors with Short Seeds

### v1

@@ 5.2.2 The (relatively) high-error case
    $Y_{\mathsf{BExt}}\circ\mathsf{BExt}(Z,Y_{\mathsf{BExt}})\approx_{\varepsilon_{\mathsf{block}}+(t+1)\varepsilon_{0}}U_{d_{t}+m_{0}}.$
We now define our final strong extractor $\mathsf{Ext}\colon\{0,1\}^{n}\times\{0,1\}^{d}\to\{0,1\}^{m_{0}}$ (recall that we abbreviate $m_{0}=m_{0}(\ell_{0})$). Choose our overall seed to be $Y=Y_{\mathsf{Cond}}\circ Y_{0}\circ\cdots\circ Y_{t}\circ Y_{\mathsf{BExt}}$ and set $\mathsf{Ext}(X,Y)=\mathsf{BExt}(Z,Y_{\mathsf{BExt}})$. By the discussion above, $\mathsf{Ext}$ is a strong $(k,\varepsilon)$-extractor with error (recall that we abbreviate $\varepsilon_{0}=\varepsilon_{0}(\ell_{0})$)
$\varepsilon=2\varepsilon_{0}+\varepsilon_{\mathsf{block}}+(t+1)\varepsilon_{0}\leq(2t+4)(\varepsilon_{0}+2\cdot 2^{-c_{\mathsf{Samp}}k})$
for a sufficiently large constant $C$ since $t=O(\log u)$ (where one of the $\varepsilon_{0}$ terms comes from fixing the seed in the condensing step of Item 1), and seed length
$d=|Y_{\mathsf{Cond}}|+|Y_{\mathsf{BExt}}|+\sum_{i=0}^{t}|Y_{i}|\leq C_{\mathsf{Cond}}\log(n/\varepsilon_{0})+d_{t}+(t+1)C_{\mathsf{Samp}}\log(n^{\prime}/\varepsilon_{0})\leq C\log u\cdot\log(n/\varepsilon_{0})$
provided that $C$ is large enough (again since $t=O(\log u)$), as desired. We used Equation 8 to bound $d_{t}$ and obtain the last inequality.
@@
It remains to analyze the time complexity of $\mathsf{Ext}$. If $k\geq C\log^{2}(n/\varepsilon_{0})$ with $C$ a sufficiently large constant, then Item 1 takes time $\widetilde{O}(n)$. Item 2 takes time $t\cdot\widetilde{O}(n)=\widetilde{O}(n)$, since $t=O(\log u)=O(\log n)$ and each averaging sampler $\mathsf{Samp}_{i}$ is computable in time $\widetilde{O}(n)$. Item 3 takes time $t\cdot\widetilde{O}(n)=\widetilde{O}(n)$, since $\mathsf{Ext}_{0}$ and each $\mathsf{Ext}_{i}$ from Lemma 2.14 are computable in time $\widetilde{O}(n)$. Therefore, $\mathsf{Ext}$ is computable in overall time $\widetilde{O}(n)$ in this case.
Otherwise, if $k<C\log^{2}(n/\varepsilon_{0})$, then Item 1 takes time $\widetilde{O}(n)$ after a preprocessing step corresponding to finding a primitive element of $\mathds{F}_{q}$ with $q\leq\operatorname{poly}(n/\varepsilon_{0})$. As discussed above, Item 2 takes time $\widetilde{O}(n)$. Item 3 takes time $\widetilde{O}(\ell_{0})=\widetilde{O}(n)$ after a preprocessing step, and so $\mathsf{Ext}$ is computable in overall time $\widetilde{O}(n)$ after a preprocessing step. Moreover, if the preprocessing step for $\mathsf{Ext}_{0}$ consists in finding primitive elements of $j$ fields $\mathds{F}_{q}$ with orders $q\leq\operatorname{poly}(n/\varepsilon_{0})$, then by the above the preprocessing step for $\mathsf{Ext}$ consists in finding primitive elements of $j+1$ fields $\mathds{F}_{q}$ with orders $q\leq\operatorname{poly}(n/\varepsilon_{0})$. ∎
@@
    if $k\geq 2^{C\cdot i}\cdot\log^{2}(n/\varepsilon)$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$;
@@
    if $k<2^{C\cdot i}\cdot\log^{2}(n/\varepsilon)$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$ after a preprocessing step which corresponds to finding primitive elements of $i$ fields $\mathds{F}_{q}$ of orders $q\leq\operatorname{poly}(n/\varepsilon)$ powers of $2$.
Consequently, if we choose $i$ to be the largest integer such that $\log^{(i)}\!n\geq 6C$ (which satisfies $i\leq\log^{*}\!n$) we get a strong $(k,\varepsilon)$-seeded extractor $\mathsf{Ext}\colon\{0,1\}^{n}\times\{0,1\}^{d}\to\{0,1\}^{m}$ with seed length $d\leq 6C^{2}\log(n/\varepsilon)$ and output length $m\geq k/2^{C\log^{*}\!n}$ for any error $\varepsilon\geq n^{3}\cdot 2^{-k/2^{C\log^{*}\!n}}$. If $k\geq 2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$. Otherwise, $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$ after a preprocessing step which corresponds to finding primitive elements of $i\leq\log^{*}\!n$ fields $\mathds{F}_{q}$ of orders $q\leq\operatorname{poly}(n/\varepsilon)$.
@@
We iteratively apply Lemma 5.9 $i$ times. Let $c,C>0$ be the constants guaranteed by Lemma 5.9. For the first application of the lemma, we take $\mathsf{Ext}_{0}\colon\{0,1\}^{n}\times\{0,1\}^{d_{0}}\to\{0,1\}^{m_{0}}$ to be the strong $(k_{0}=0.7n,\varepsilon_{0})$ extractor from Lemma 2.14 with $m_{0}=k_{0}/20$ and $\varepsilon_{0}\geq 2^{-ck_{0}/100}$ to be defined later. The corresponding seed length is $d_{0}\leq 2m_{0}+4\log(n/\varepsilon_{0})+4$, which satisfies $d_{0}\leq k_{0}$, and so the initial value of $u$ is $u_{0}=d_{0}/\log(n/\varepsilon_{0})\leq k_{0}$. Denote by $\mathsf{Ext}_{1}$ the resulting strong seeded extractor. In the second application of Lemma 5.9, we instantiate $\mathsf{Ext}_{0}$ with $\mathsf{Ext}_{1}$ instead to obtain a new strong seeded extractor $\mathsf{Ext}_{2}$, and so on. For each $j\in[i]$, we obtain a family of strong $(k,\varepsilon_{j})$-seeded extractors $\mathsf{Ext}_{j}\colon\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d_{j}}\to\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{m_{j}}$ parameterized by $k$ with output length $m_{j}=m_{j-1}(ck)$, error
@@
Finally, we discuss the time complexity of $\mathsf{Ext}$. Note that the initial choice for $\mathsf{Ext}_{0}$ is computable in time $\widetilde{O}(n_{0})$. Therefore, if $k\geq 2^{C\cdot i}\log^{2}(n/\varepsilon)$ for a sufficiently large constant $C>0$, then the conditions of Item 1 of Lemma 5.9 are satisfied for all $i$ applications of this lemma, and so $\mathsf{Ext}$ will be computable in time $\widetilde{O}(n)$. Otherwise, the condition in Item 1 holds and so $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$ after a preprocessing step, since we always have $u\leq n$ in each application of the lemma. By Lemma 5.9, the preprocessing amounts to finding primitive elements of $i$ fields $\mathds{F}_{q}$ with orders $q\leq\operatorname{poly}(n/\varepsilon_{0})=\operatorname{poly}(n/\varepsilon)$. ∎
To obtain our final theorem, we use block source extraction to increase the output length of the extractor from Corollary 5.10, following a strategy of Zuckerman [Zuc97].
@@
    if $k\geq 2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$;
@@
    if $k<2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$ after a preprocessing step which corresponds to finding $\log^{*}n$ primitive elements of fields $\mathds{F}_{q}$ of orders $q\leq\operatorname{poly}(n/\varepsilon)$ powers of $2$.
@@
Define $\varepsilon^{\prime}=\varepsilon/6$ and let $X$ be an arbitrary $(n,k)$-source. The extractor $\mathsf{Ext}$ behaves as follows on input $X$:
@@
    Apply a strong $(k,k^{\prime},(\varepsilon^{\prime})^{2})$-condenser $\mathsf{Cond}\colon\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d_{\mathsf{Cond}}}\to\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n^{\prime}}$ to $X$, with output min-entropy rate $k^{\prime}\geq 0.95n^{\prime}$ and seed length $d_{\mathsf{Cond}}=C_{\mathsf{Cond}}\log(n/\varepsilon^{\prime})$. If $k\geq 2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$, we instantiate $\mathsf{Cond}$ with the lossless KT strong $(k,k^{\prime},\varepsilon^{\prime})$-condenser (Theorem 3.10). Otherwise, we instantiate $\mathsf{Cond}$ with the lossy RS strong $(k,k^{\prime},\varepsilon^{\prime})$-condenser (Theorem 3.11). By the second part of either Theorem 3.10 or Theorem 3.11, we get that with probability at least $1-\varepsilon^{\prime}$ over the choice of the seed $y$ we obtain an output $X^{\prime}$ that is $\varepsilon^{\prime}$-close to an $(n^{\prime},k^{\prime})$-source with $k^{\prime}\geq 0.95n^{\prime}$. As in previous arguments, we work under such a good fixing of $y$ from here onwards and account for the probability $\varepsilon^{\prime}$ of selecting a bad seed in the final extractor error later on.
@@
    Write $X^{\prime}=X_{1}\circ X_{2}$ with $|X_{1}|=|X_{2}|=n^{\prime}/2$. Choose the constant $c>0$ in the theorem statement small enough so that $\log(1/\varepsilon^{\prime})\leq\log(1/\varepsilon)+3\leq ck+3\leq 0.05k$, which means that $n^{\prime}/2-0.05k-\log(1/\varepsilon^{\prime})\geq 0.4n^{\prime}$. Then, combining Item 1 with Lemma 2.24, (instantiated with $t=2$, $\Delta=0.05k$, and $\varepsilon=\varepsilon^{\prime}$) via the triangle inequality, $X^{\prime}$ is $3\varepsilon^{\prime}$-close to an $((n^{\prime}/2,n^{\prime}/2),0.8)$-block source.
@@
    Apply block source extraction to $X_{1}\circ X_{2}$. More precisely, let $\mathsf{Ext}_{1}\colon\{0,1\}^{n_{1}}\times\{0,1\}^{d_{1}}\to\{0,1\}^{m_{1}}$ be the strong $(k_{1}=0.8n_{1},\varepsilon_{1})$-seeded extractor from Corollary 5.10 instantiated with $i=2$ and $n_{1}=n^{\prime}/2$, which yields $\varepsilon_{1}=\varepsilon\geq n_{1}^{3}\cdot 2^{-c_{1}k_{1}}$, $d_{1}\leq C_{1}\log\log k_{1}\cdot\log(n^{\prime}/\varepsilon)$, and $m_{1}\geq c_{1}k_{1}$, for constants $c_{1},C_{1}>0$ guaranteed by Corollary 5.10. Furthermore, let $\mathsf{Ext}_{2}\colon\{0,1\}^{n_{2}}\times\{0,1\}^{d_{2}}\to\{0,1\}^{m_{2}}$ be the strong $(k_{2}=0.8n_{2},\varepsilon_{2})$-seeded extractor from the “Consequently” part of Corollary 5.10 and $n_{2}=n^{\prime}/2$, which yields $\varepsilon_{2}=n_{2}^{3}\cdot 2^{-k_{2}/2^{C_{2}\log^{*}\!k_{2}}}$, $d_{2}\leq C_{2}\log(n^{\prime}/\varepsilon)$, and $m_{2}\geq k_{2}/2^{C_{2}\log^{*}\!k_{2}}$, for a constant $C_{2}>0$ guaranteed by Corollary 5.10. This choice of parameters ensures that $m_{2}\geq d_{1}$. Indeed, since $k\geq k_{1}=k_{2}\geq 0.4n^{\prime}$, to see that $m_{2}\geq d_{1}$ it suffices to check that
@@
    Since $\varepsilon_{1}=\varepsilon^{\prime}=\varepsilon/5$ and $\log(n^{\prime}/\varepsilon_{1})=O(\log(k/\varepsilon^{\prime}))=O(\log k+k/\log k)=O(k/\log k)$, it is enough that

### v2

@@ 4.2.2 The (relatively) high-error case
    $\mathopen{}\mathclose{{\left(Y_{0},\dots,Y_{t},Y_{\mathsf{BExt}},\mathsf{BExt}(Z,Y_{\mathsf{BExt}})}}\right)\approx_{\varepsilon_{\mathsf{block}}+(t+1)\varepsilon_{0}}\mathopen{}\mathclose{{\left(Y_{0},\dots,Y_{t},Y_{\mathsf{BExt}},U_{m_{0}}}}\right).$
We now define our final strong extractor $\mathsf{Ext}\colon\{0,1\}^{n}\times\{0,1\}^{d}\to\{0,1\}^{m_{0}}$ (recall that we abbreviate $m_{0}=m_{0}(\ell_{0})$). Choose our overall seed to be $Y=(Y_{\mathsf{Cond}},Y_{0},Y_{1},\dots,Y_{t},Y_{\mathsf{BExt}})$ and set $\mathsf{Ext}(X,Y)=\mathsf{BExt}(Z,Y_{\mathsf{BExt}})$. By the discussion above, $\mathsf{Ext}$ is a strong $(k,\varepsilon)$-extractor with error (recall that we abbreviate $\varepsilon_{0}=\varepsilon_{0}(\ell_{0})$)
$\varepsilon=2\varepsilon_{0}+\varepsilon_{\mathsf{block}}+(t+1)\varepsilon_{0}\leq(2t+4)(\varepsilon_{0}+2\cdot 2^{-\beta_{\mathsf{Samp}}k}).$
As discussed above, the $2\varepsilon_{0}$ accounts for fixing the seed in the condensing step of item˜1 and for assuming that $X^{\prime}$ is an $(n^{\prime},k^{\prime})$-source under this fixing. Since $t=O(\log u)$, if we pick $C>0$ to be a sufficiently large constant and $c$ to be smaller than $\beta_{\mathsf{Samp}}$ so that $\varepsilon_{0}\geq 2^{-ck_{0}}\geq 2^{-ck}\geq 2^{-\beta_{\mathsf{Samp}}k}$, we get
$\varepsilon\leq C\log u\cdot\varepsilon_{0}.$
The seed length is
$d=|Y_{\mathsf{Cond}}|+|Y_{\mathsf{BExt}}|+\sum_{i=0}^{t}|Y_{i}|\leq C_{\mathsf{Cond}}\log(n/\varepsilon_{0})+d_{t}+t\cdot\gamma\log(n^{\prime}/\varepsilon_{0})\leq C\log u\cdot\log(n/\varepsilon_{0}),$
where $\gamma=\max(C^{\prime}_{\alpha},\gamma_{\mathsf{Samp}})$, provided that $C$ is large enough (again since $t=O(\log u)$), as desired. We used eq.˜19 to bound $d_{t}$ and obtain the last inequality.
@@
It remains to analyze the time complexity of $\mathsf{Ext}$. We proceed by cases:
  • If $k\geq C\log^{2}(n/\varepsilon_{0})$ with $C$ a sufficiently large constant, then by theorem˜3.14 we get that item˜1 either takes time $\widetilde{O}(n+\sqrt{n}\cdot\log(1/\varepsilon_{0})^{5})$, or time $\widetilde{O}(n)$ after a one-time preprocessing step. Regarding item˜2, each averaging sampler $\mathsf{Samp}_{i}$ from lemma˜2.20 runs in time $\log^{2}(1/\varepsilon_{0})\cdot\operatorname{polylog}n+O(\ell_{i}\log n)$. This is $\widetilde{O}(n)$ when $\varepsilon_{0}\geq 2^{-\widetilde{O}(\sqrt{n})}$, which is implied by the constraint $k\geq C\log^{2}(n/\varepsilon_{0})$. Since there are $t=O(\log u)=O(\log n)$ blocks, item˜2 runs in $\widetilde{O}(n)$ times. item˜3 takes time $T(\ell_{0})+t\cdot\widetilde{O}(n)=T(\ell_{0})+\widetilde{O}(n)\leq T(n)+\widetilde{O}(n)$, since $\mathsf{Ext}_{0}$ is computable in time $T(\ell_{0})$, each $\mathsf{Ext}_{i}$ from lemma˜2.13 are computable in time $\widetilde{O}(n)$, and $\ell_{0}\leq n$. Therefore, in this case $\mathsf{Ext}$ is computable in overall time $T(n)+\widetilde{O}(n+\sqrt{n}\cdot\log(1/\varepsilon_{0})^{5})$, or time $T(n)+\widetilde{O}(n)$ after a preprocessing step.
  • If $k<C\log^{2}(n/\varepsilon_{0})$, then item˜1 takes time $\widetilde{O}(n)$ after a preprocessing step. In this case, item˜2 amounts to $t=O(\log n)$ applications of theorem˜3.17, and so runs in time $\widetilde{O}(n)$ after a preprocessing step. item˜3 takes time $T(\ell_{0})+\widetilde{O}(n)$ after a preprocessing step, and so $\mathsf{Ext}$ is computable in overall time $T(\ell_{0})+\widetilde{O}(n)\leq T(n)+\widetilde{O}(n)$ after a preprocessing step.
@@
    if $k\geq 2^{C\cdot i}\cdot\log^{2}(n/\varepsilon)$ and $\varepsilon\geq 2^{-Cn^{0.1}}$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$;
@@
    if $k<2^{C\cdot i}\cdot\log^{2}(n/\varepsilon)$ or $\varepsilon<2^{-Cn^{0.1}}$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$ after a preprocessing step.
Consequently, if we choose $i$ to be the largest integer such that $\log^{(i)}\!n\geq 6C$ (which satisfies $i\leq\log^{*}\!n$) we get a strong $(k,\varepsilon)$-seeded extractor $\mathsf{Ext}\colon\{0,1\}^{n}\times\{0,1\}^{d}\to\{0,1\}^{m}$ with seed length $d\leq 6C^{2}\log(n/\varepsilon)$ and output length $m\geq k/2^{C\log^{*}\!n}$ for any error $\varepsilon\geq n^{3}\cdot 2^{-k/2^{C\log^{*}\!n}}$. If $k\geq 2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$ and $\varepsilon\geq n^{3}\cdot 2^{-Cn^{0.1}}$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$. Otherwise, $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$ after a preprocessing step.
@@
This is a consequence of iteratively applying lemma˜4.9 $i$ times. Note that here part of the relevant condition for the preprocessing is $k\geq 2^{Ci}\log^{2}(n/\varepsilon)$. The reason behind this is that each application of lemma˜4.9 reduces the min-entropy requirement by a constant factor.
Let $c,C>0$ be the constants guaranteed by lemma˜4.9. For the first application of the lemma, we take $\mathsf{Ext}_{0}\colon\{0,1\}^{n}\times\{0,1\}^{d_{0}}\to\{0,1\}^{m_{0}}$ to be the strong $(k_{0}=0.7n,\varepsilon_{0})$ extractor from lemma˜2.13 with $m_{0}=k_{0}/20$ and $\varepsilon_{0}\geq 2^{-ck_{0}/100}$ to be defined later. The corresponding seed length is $d_{0}\leq 2m_{0}+4\log(n/\varepsilon_{0})+4$, which satisfies $d_{0}\leq k_{0}$, and so the initial value of $u$ is $u_{0}=d_{0}/\log(n/\varepsilon_{0})\leq k_{0}$. Denote by $\mathsf{Ext}_{1}$ the resulting strong seeded extractor. In the second application of lemma˜4.9, we instantiate $\mathsf{Ext}_{0}$ with $\mathsf{Ext}_{1}$ instead to obtain a new strong seeded extractor $\mathsf{Ext}_{2}$, and so on. For each $j\in[i]$, we obtain a family of strong $(k,\varepsilon_{j})$-seeded extractors $\mathsf{Ext}_{j}\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}\times\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d_{j}}\to\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{m_{j}}$ parameterized by $k$ with output length $m_{j}=m_{j-1}(ck)$, error
@@
Finally, we discuss the time complexity of $\mathsf{Ext}$. Note that the initial choice for $\mathsf{Ext}_{0}$ is computable in time $\widetilde{O}(n_{0})$. Therefore, if $k\geq 2^{C\cdot i}\log^{2}(n/\varepsilon)$ then each application of lemma˜4.9 runs in time $\widetilde{O}(n+\sqrt{n}\log(1/\varepsilon_{0}(c^{i}k))^{5})$. This uses the fact that the error increases in each application of lemma˜4.9. Since $\varepsilon_{0}(c^{i}k)\geq\varepsilon/n^{3}$, then each application of lemma˜4.9 runs in time $\widetilde{O}(n)$ without preprocessing when $\varepsilon\geq 2^{-Cn^{0.1}}$. Otherwise, the condition in footnote˜27 of lemma˜4.9 holds and so $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$ after a preprocessing step, since we always have $u\leq n$ in each application of the lemma. ∎
Remark 4.11 (the preprocessing in corollary˜4.10).
For the sake of readability we were not explicit about the precise preprocessing in the statement of Case 2 of corollary˜4.10. We expand on that now. corollary˜4.10 recursively invokes lemma˜4.9 $i$ times. All these recursive calls use the same type of condenser (either the KT condenser or the RS condenser), depending on the initial choices of $k$ and $\varepsilon$. Therefore:
  • If $k\geq 2^{C\cdot i}\cdot\log^{2}(n/\varepsilon)$ and $\varepsilon<2^{-Cn^{0.1}}$, then the preprocessing corresponds to the preprocessing for $i$ calls of the KT condenser – for example, generating $i$ primes $q\leq\operatorname{poly}(n/\varepsilon)$.
  • If $k<2^{C\cdot i}\cdot\log^{2}(n/\varepsilon)$ and $\varepsilon<2^{-Cn^{0.1}}$, then each invocation of lemma˜4.9 requires one preprocessing step for the RS condenser call in item˜1, and, naively, $t$ preprocessing steps for the $t=O(\log n)$ RS condenser calls in item˜2b. We can further improve this by noting that all $t$ calls of theorem˜3.17 in item˜2b use the same input length, error, and $\alpha$, and so we only need to run one preprocessing step that suffices for all $t$ calls simultaneously. So, overall, the preprocessing corresponds to the preprocessing for $2i$ calls of the RS condenser – generating $2i$ primes $q\leq\operatorname{poly}(n/\varepsilon)$ along with a primitive element for each $\mathds{F}_{q}$.
Note that corollary˜4.10 only guarantees output length $k/2^{Ci}$ for each $i$. In particular, when $i=\log^{*}\!n$ we get output length $k/2^{C\log^{*}\!n}$. This is slightly sublinear, and we would like to aim for output length $ck$ for some constant $c>0$. To obtain our final theorem, we use block source extraction to increase the output length of the extractor from corollary˜4.10, following a strategy of Zuckerman [ZUC97].
@@
    if $k\geq 2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$ and $\varepsilon\geq 2^{-Cn^{0.1}}$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$;
@@
    if $k<2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$ or $\varepsilon<2^{-Cn^{0.1}}$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$ after a preprocessing step.
@@
We begin by providing an informal discussion of the proof. On input an arbitrary $(n,k)$-source $X$, we begin by applying a fast condenser to $X$ to obtain another source $X^{\prime}$ with high min-entropy rate. Then, we split $X^{\prime}$ in half to obtain a block source $(X_{1},X_{2})$ with two blocks. Then, we perform block source extraction on $(X_{1},X_{2})$. More concretely, we apply the extractor obtained by instantiating corollary˜4.10 with $i=\log^{*}\!n$ to $X_{2}$, and then use its output as the seed to extract from $X_{1}$ using the extractor obtained by instantiating corollary˜4.10 with $i=2$ (which has output length $\Omega(k)$).
More formally, define $\varepsilon^{\prime}=\varepsilon/6$ and let $X$ be an arbitrary $(n,k)$-source. The extractor $\mathsf{Ext}$ behaves as follows on input $X$:
@@
    Apply a strong $(k,k^{\prime},(\varepsilon^{\prime})^{2})$-condenser $\mathsf{Cond}\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}\times\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d_{\mathsf{Cond}}}\to\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n^{\prime}}$ to $X$, with output min-entropy $k^{\prime}\geq 0.95n^{\prime}$ and seed length $d_{\mathsf{Cond}}=C_{\mathsf{Cond}}\log(n/\varepsilon^{\prime})$. If $k\geq 2^{C\log^{*}\!n}\cdot\log^{2}(n/\varepsilon)$, we instantiate $\mathsf{Cond}$ with the KT strong $(k,k^{\prime},\varepsilon^{\prime})$-condenser (theorem˜3.14 instantiated with $\alpha=0.05$). Otherwise, we instantiate $\mathsf{Cond}$ with the RS strong $(k,k^{\prime},\varepsilon^{\prime})$-condenser (theorem˜3.17 instantiated with $\alpha=0.025$). By the second part of either theorem˜3.14 or theorem˜3.17, we get that with probability at least $1-\varepsilon^{\prime}$ over the choice of the seed $y$ we obtain an output $X^{\prime}$ that is $\varepsilon^{\prime}$-close to an $(n^{\prime},k^{\prime})$-source with $k^{\prime}\geq 0.95n^{\prime}$. As in previous arguments, we work under such a good fixing of $y$ from here onwards and account for the probability $\varepsilon^{\prime}$ of selecting a bad seed in the final extractor error later on.
@@
    Let $(X_{1},X_{2})$ correspond to the first two blocks of $\lfloor n^{\prime}/2\rfloor$ bits of $X^{\prime}$. Choose the constant $c>0$ in the theorem statement small enough so that $\log(1/\varepsilon^{\prime})\leq\log(1/\varepsilon)+3\leq ck+3\leq 0.05k-1$, which means that $\lfloor n^{\prime}/2\rfloor-0.05k-\log(1/\varepsilon^{\prime})\geq 0.4n^{\prime}$. Then, combining item˜1 with lemma˜2.23 (instantiated with $t=2$, $\Delta=0.05k$, and $\varepsilon=\varepsilon^{\prime}$) via the triangle inequality, $(X_{1},X_{2})$ is $3\varepsilon^{\prime}$-close to an exact $((n_{1}=\lfloor n^{\prime}/2\rfloor,n_{2}=\lfloor n^{\prime}/2\rfloor),0.8)$-block-source.
@@
    Apply block source extraction to $(X_{1},X_{2})$. More precisely, let $\mathsf{Ext}_{1}\colon\{0,1\}^{n_{1}}\times\{0,1\}^{d_{1}}\to\{0,1\}^{m_{1}}$ be the strong $(k_{1}=0.8n_{1},\varepsilon_{1}=\varepsilon^{\prime})$-seeded extractor from corollary˜4.10 instantiated with $i=2$ and $n_{1}=\lfloor n^{\prime}/2\rfloor$, which requires $\varepsilon_{1}=\varepsilon^{\prime}\geq n_{1}^{3}\cdot 2^{-c_{1}k_{1}}$ and guarantees $d_{1}\leq C_{1}\log\log k_{1}\cdot\log(n^{\prime}/\varepsilon)$ and $m_{1}\geq c_{1}k_{1}$, for constants $c_{1},C_{1}>0$ guaranteed by corollary˜4.10. Furthermore, let $\mathsf{Ext}_{2}\colon\{0,1\}^{n_{2}}\times\{0,1\}^{d_{2}}\to\{0,1\}^{m_{2}}$ be the strong $(k_{2}=0.8n_{2},\varepsilon_{2}=\varepsilon^{\prime})$-seeded extractor from the “Consequently” part of corollary˜4.10 and $n_{2}=\lfloor n^{\prime}/2\rfloor$, which requires $\varepsilon_{2}\geq n_{2}^{3}\cdot 2^{-k_{2}/2^{C_{2}\log^{*}\!k_{2}}}$ and guarantees $d_{2}\leq C_{2}\log(n^{\prime}/\varepsilon)$ and $m_{2}\geq k_{2}/2^{C_{2}\log^{*}\!k_{2}}$, for a constant $C_{2}>0$ guaranteed by corollary˜4.10. This choice of parameters ensures that $m_{2}\geq d_{1}$ and is valid by the lower bound on $\varepsilon$ in the theorem statement, recalling that $\varepsilon^{\prime}=\varepsilon/6$. Indeed, since $k\geq k_{1}=k_{2}\geq 0.4n^{\prime}$, to see that $m_{2}\geq d_{1}$ it suffices to check that
@@
    Since $\varepsilon_{1}=\varepsilon^{\prime}=\varepsilon/6$ and $\log(n^{\prime}/\varepsilon_{1})=O(\log(k/\varepsilon^{\prime}))=O(\log k+k/\log k)=O(k/\log k)$, it is enough that